Following information is available in respect of a company named, as Beta Ltd
Particulars Amount
Profit before taxes Rs 50 lakhs
Dividend declared per share Rs2
Price of the share prevalent on stock
exchange Rs 200
Applicable tax rate 35%
Share capital of company 5 lakh shares
of Rs 10 each

Calculate-
 Earnings per share
 Dividend Pay-out Ratio
 Price earnings ratio

b. Prepare the cash flow statement from investing activities of Alpha Creative Ltd for the
year ended March31, 2019
1
Expert's answer
2019-10-09T10:54:29-0400

a. Earnings per share = Profit/Share capital,


"EPS = \frac{50×(1 - 0.35)lakhs}{5lakhs} = 6.5."

Dividend Pay-out Ratio = Dividend per share/Earnings per share,


"DP ratio = \frac{2}{6.5} = 0.31."

Price earnings ratio = Price per share/Earnings per share,


"PE ratio = \frac{200}{6.5} = 30.77."

b. Cash flow from investing activities is the section of a company's cash flow statement that displays how much money has been used in (or generated from) making investments during a specific time period.